COUPLING OF NON-SYMMETRIC BENDING MODES TO METASTABLE STATES OF ACETYLENE

\underline{KYLE L. BITTINGER}, WILTON L. VIRGO, and ROBERT W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139.

Several bands of acetylene involving the near degenerate non-symmetric bending modes ν_{4} and ν_{6} of S_{1} have been recorded by simultaneous laser induced fluorescence and gold surface electron ejection by laser-excited metastables (Au:SEELEM). The Au:SEELEM spectra indicate pronounced differences in coupling to long-lived states between the torsion (ν_{4}) and in-plane bend (ν_{6}). These results will be examined in light of the fact that ab initio calculations predict a nonplanar equilibrium geometry for the third triplet state, which mediates coupling between S_{1} and the triplet manifold.